For each of the following distances, if a galaxy the size of the Milky Way (about 30.1 kpc in diameter) were at that distance from the Milky Way, what would its angular size appear to be?

11.7 Mpc - a = --- arc minutes

1034 Mpc- a = ---arc mintues
Answer #1

Angular Size is also used as a measure of optical instruments resolving power - basically how small an object can be seen. Angular Size refers to the object's apparent size as seen from an observer on Earth.

The galaxy size of the milky way is 30.1 kpc in diameter

To get the diameter in kpc

D = 1000 * d (2\pi A) /(360*3600)

D = d(\piA) / 648

D = d A / 206

where d is in Mpc  , D is in kpc and A is arcsec

1) d = 11.7 mpc, D = 30.1 kpc

30.1= (11.7) * (A) / 206

30.1 *206 = 11.7 *A

A = 529.96 arcsec

2) d = 1034 mpc, D = 30.1 kpc

30.1 = (1034 * A ) / 206

A = 5.99 arc sec